Binomial Distribution. Research shows that in the U.S. federal courts, about 90% of defendants are found guilty in criminal trials. Suppose we take a random sample of 25 trials. (For this problem it is best to use the Binomial Tables).

Based on a proportion of .90, what is the expected value( mean) distribution?
Answer #1

Solution :

Given that,

p = 90%=0.90

q = 1 - p =1-0.90=0.10

n = 25

Using binomial distribution,

\mu = n * p = 25*0.90=22.5